OpenStudy (anonymous):

how do you solve in scientific notation such as the 25.63 in scientific notation

thomaster (thomaster):

Move the point to the left till there's only 1 number before the point. You begin with \(\large 25.63\times 10^0\) With every movement of the point to the left, you add 1 to the exponent after 10. so 1 to the left and it'll become \(\large 2.563\times 10^1\) Now you usually only use 2 significant figures AFTER the point so it will be \(\large 2.56\times 10^1\)

OpenStudy (anonymous):

thanks @thomaster